### v3

# Swarm-intelligent search for gravitational waves from eccentric binary mergers

Abstract
We implement an eccentric search for compact binary mergers based on particle swarm optimization. Orbital eccentricity is an invaluable input for understanding the formation scenarios of the binary mergers and can play a pivotal role in finding their electromagnetic counterparts. Current modelled searches rely on pre-computed template banks that are computationally expensive and resistant towards expanding the search parameter space dimensionality. On the other hand, particle swarm optimization offers a straightforward algorithm that dynamically selects template points while exploring an arbitrary dimensional parameter space. Through extensive evaluation using simulated signals from spin-aligned eccentric binary mergers, we discovered that the search exhibits a remarkable autonomy in capturing the effects of both eccentricity and spin. We describe our search pipeline and revisit some of the merger candidates from the gravitational wave transient catalogs.
I Introduction
Till the third observing run (O3), the Advanced LIGOs and the Advanced Virgo have collectively contributed to over 90 confident detections of transient gravitational wave (GW) events (Abbott et al., 2019a, 2021a, 2021b, 2021c). These include the first observations of binary black holes (BBHs), binary neutron stars (BNSs), and the neutron star-black hole (NSBH) systems (Abbott et al., 2016, 2017, 2021d). In addition, several marginal candidate events have also been reported. Geographic separation of Virgo from the twin LIGO detectors has often aided the localization of the sources in the sky. The KAGRA and the GEO600 detectors also participated but their detection capabilities were marginal (Collaboration et al., 2022). With the addition of the LIGO-India observatory later in this decade, the network of second generation detectors should be complete (Saleem et al., 2021).
Probing the orbital eccentricity of compact binary coalescences (CBCs) is vital for understanding their formation mechanism. Its accurate measurement on a sizeable population of merger events can help to decide over the dominant formation process, for example a dynamical capture or an isolated evolution. The earlier understanding was that the eccentricity could radiate away quickly, and the orbit becomes circularized before the signal enters the sensitive band of the detectors. However, it may be possible that binary systems formed via a dynamical interaction would not have enough time to circularize and can carry a non-zero eccentricity. Although the LVK Collaboration have not reported eccentric events in the past (Abbott et al., 2019b), there are several attempts to find signatures of eccentricity in the LVK events (Wu et al., 2020; Romero-Shaw et al., 2019; Gayathri et al., 2022; Romero-Shaw et al., 2021, 2022, 2020a, 2020b; Lenon et al., 2020; Bonino et al., 2023; Gamba et al., 2023; Iglesias et al., 2022). Currently, these efforts are not concerned with the detection stage but are usually conducted as offline parameter estimation with schemes based on Bayesian techniques. However, studies that infer binaries’ formation channels based on the eccentric parameter estimation of events flagged by non-eccentric searches could be biased.
Nevertheless, significant developments towards eccentric parameter estimation have been made, for example using a hybrid approach- by doing an initial analysis with a non-precessing waveform like the IMRPhenomD and then reweighing the results with the eccentricity-marginalised Bayesian likelihood using the SEOBNRE model (Romero-Shaw et al., 2019, 2020a, 2023, 2021, 2022, 2020b). Studies have indicated that effects due to orbital eccentricity and spin-induced orbital precession can be correlated (Romero-Shaw et al., 2020a, 2023; Bustillo et al., 2021). The lack of accurate and fast-to-generate eccentric waveforms describing the full set of binary parameters hinders their implementation in the LVK analyses. Developing eccentric waveforms with spins is an active area of research in the community. These waveforms are mostly intended for parameter estimation studies. Note that currently, the description of eccentricity is not standardized and can vary across different models (Knee et al., 2022; Shaikh et al., 2023). However, defining eccentricity at 10 Hz is almost a standard practice.
The standard technique to search for CBC signals in the detector data involves cross-correlating the detector output with a bank of template waveforms that attempt to model the gravitational radiation. The current searches look for the spin effects in terms of the aligned or anti-aligned spins of the merger components, in addition to the component masses (Messick et al., 2017; Adams et al., 2016; Usman et al., 2016; Chu et al., 2022). In principle, including extensive parameters in the templates describing the source should benefit the search process. In practice, however, the benefit critically hinges on the compuational cost and the feasibility of implementation. It is non-trivial to further expand the parameter dimensionality of the current template-bank based searches. However, if an inspiraling binary neutron star signal has a residual eccentricity while entering the detector sensitive band, it would modulate the signal for a large number of low frequency cycles before fading through its late inspiral phase. Thus, a significant fraction of the signal-to-noise ratio (SNR) can be accumulated from the low-frequency inspiral part, potentially improving its significance and aiding the sky localization. However, the LVK collaboration currently uses only the aligned-spin searches for generating candidates for their follow-up. Also note that BAYESTAR, which computes the skymaps in low latencies, is not optimized for localizing eccentric sources (Singer and Price, 2016).
An earlier search targeting non-spinning eccentric BNSs using a stochastic template placement algorithm could not lead to any interesting eccentric BNS candidate (Nitz et al., 2020). If eccentricity is associated with some of the detectable events, a generic eccentric search can enhance the significance of some marginal events flagged by the aligned-spin searches. After this paper was originally submitted, a search appeared for neutron star binaries with a template bank that is about two orders of magnitude larger than the aligned-spin bank (Dhurkunde and Nitz, 2023). That search reports no new significant candidate from the third Observing Run of the Advanced LIGO and Advanced Virgo.
In this work, we demonstrate a spin-aligned eccentric search using the Particle Swarm Optimization (PSO) algorithm. Previously, PSO based searches have been explored towards conducting aligned-spin and even fully precessing searches but never included eccentricity (Wang and Mohanty, 2010; Srivastava et al., 2018; Normandin and Mohanty, 2020). Here, we include the orbital eccentricity as a search parameter in addition to the component masses and their aligned-spins. To generate templates waveforms for BNS systems, we use the inspiral-only TaylorF2Ecc model. It is a fast, frequency-domain model supporting small eccentricities in the presence of aligned-spins (Moore et al., 2016). For injecting signals from simulated sources, we primarily use the eccentric version of the time-domain TEOBResumS model (called TEOBResumS-DALI) that can incorporate very high eccentricities in the presence of aligned-spins (Chiaramello and Nagar, 2020; Nagar et al., 2021). Here we also use it for generating templates for BBH systems. Currently, neither of these waveform models incorporate the mean anomaly parameter and thus, all our injections and the results are obtained at a fixed zero mean anomaly. Our key observations and insights are summarized as follows.
1.
To explore a given search parameter space, the PSO algorithm iteratively computes the templates points on the fly while optimizing a given detection statistic. This approach is different from that of a template-bank which depends on a fixed, pre-determined set of template points. The PSO algorithm thus, avoids the minimal-match criterion implemented in the template-bank searches but spends more time in exploring the region of the parameter space where the signal is more likely to be found. Some parts of the search parameter space generally require larger sampling than others. For example, the signal recovery in the low mass BNS region need more template points than that in the heavier BBHs. When data from more than one detector is available, the PSO algorithm can optimize a statistic computed over the detector network.
2.
Matched-filtering of the data is preceded by template waveform generation in the PSO framework. Eccentric templates are in general costlier than the quasi-circular ones impacting the overall cost of the search. We find that the above strategy picks the aligned-spin and the eccentricity independently from each other. However, we are currently limited by the parameter space coverage of the waveform models and the speed of the dynamic template generation.
3.
Eccentricity primarily modulates the inspiral component of the signals and the gain is thus, from the low frequency content of the signals. On the other end, truncating signals till Hz by downsampling reduces the cost of matched-filtering with little to no impact on recovered the SNRs. If eccentricity is absent in the signals, the eccentric search approximately behaves like the usual aligned-spin search in terms of signal recovery.
This article is organized as follows. Section II, outlines the general search mechanism. First we describe the standard PSO algorithm in locating the optimal template for a given segment of data. We then implement a coincident search algorithm for finding event candidates with multiple detetors. Finally we discuss how a statistical significance is assigned to the coincident events. Section III demonstrates its effectiveness on simulated and real sources. We show the accuracy of recovering the injected parameters and the SNRs using some agnostic populations of BNS and BBH systems from coloured Gaussian noise. We discuss some challenges in vetoing the noise artefacts from the real instruments. We also report the re-analysis of a majority of gravitational wave transient catalog (GWTC) events with the proposed eccentric search. We finally include a parallelization scheme to reduce the runtime of the search in section III.5.
II Search Algorithm
II.1 Detection principle
In this section, we briefly review the overall method for coincident detection of compact binary coalescences. The data recorded by a gravitational wave detector is a timeseries of discretely sampled strain, containing occasional gravitational signals in the detector’s noise. The detection problem is to essentially identify instances of transient astrophysical signals in the timeseries data. This is achieved by cross-correlating the strain data with a modeled form of the astrophysical signal, called a template waveform, . A template waveform depends on several parameters, modelling the coalescing binary. The cross-correlation is performed in the frequency domain where the noise power spectral density, is used to down-weight the effect of noise, in a process called matched-filtering. The output of matched-filtering is the signal-to-noise ratio, defined as,
where a tilde represents a function in the frequency domain. Whenever there is a close match between the true signal in the data and a template waveform, the SNR timeseries attains a peak at the arrival time of the signal in the detector as shown in the Fig. 1. However at times, detector noise can also mimic GW signals by causing sharp peaks in the SNR timeseries. These occur mainly due to the non-stationarity and non-Gaussian nature of the detector noise and are called glitches. Glitches increase the chances of false alarms and affect the sensitivity to true astrophysical signals. The peaks in the SNR timeseries above a chosen threshold which arise due to the noise glitches and any potential astrophysical signals are considered as triggers. These undergo several consistency checks to veto out possible glitches. To get rid of some of the obvious loud noise transients, the signal-based power discriminator is used (Allen, 2005). Other vetoing techniques are also available for the noise triggers that dodge the standard consistency checks, especially from short duration templates (Nitz, 2018). The discriminators are designed to down-rank noise triggers in a quantity called the reweighted SNR while preserving the original SNR values for the triggers with astrophysical origin (Usman et al., 2016). Remaining triggers are tested for time-coincidence with multiple detectors - the arrival times of astrophysical signals at any pair of detectors cannot differ by more than the time-of-flight for light between the detectors with some tolerance for the measurement errors. Events that qualify these tests are assigned a statistical significance based on the rate of false alarms, which is described in section II.4. The general detection problem is extensively discussed in the literature, for example (Sathyaprakash and Schutz, 2009; Schutz, 1997; Sathyaprakash and Dhurandhar, 1991).
II.2 Swarm-intelligence
Identifying a potential signal in a given segment of data hinges on finding a close match between the signal and a template waveform. As described in the previous section, one needs to optimize function as given in the Eq. (1) over the desired range of parameters . In this section, we describe how the proposed algorithm locates the optimal template position which is represented by . The algorithm begins with a fixed number of randomly distributed template points which we refer to as a swarm. The individual members of the swarm, which are known as particles, communicate and share the knowledge gained during the search process. Each particle with a position and a velocity , where is the particle identifier and is the dimension index, constitutes a candidate for the optimal template . The initial random distribution of the particles can be chosen to be uniform in each of the search dimensions. The role of the algorithm is to iteratively improve a given fitness function by proposing a new set of positions to the particles and computing the fitness function at the newly proposed locations. The swarm eventually converges over few tens of iterations (typically less than 20) while the value of the fitness function saturates at an optimal value. Note that the fitness function must be carefully chosen since its repeated evaluation can be costly. For example, the simplest choice is the matched-filter SNR maximized over time for a single detector with Gaussian noise. However, a noise glitch can cause a very large SNR value and can deceit the algorithm. Thus we replace the SNR with the reweighted SNR for the network of detectors for maximization, as discussed in Section-II.3. To take any subsequent step, a particle considers the following three factors-
the inertial factor - the tendency to continue in the same direction as in the previous step,
the cognitive factor - the homesickness to go back to the best location achieved by itself, and
the social factor - the longing to join the best pool of peer achievers.
So at any given step, the algorithm identifies two quantities: (a) the template point with the largest fitness function among all the particles, which is called global-best or g-best, denoted as , and (b) the template point with the highest fitness achieved by the particle till the current step, which is called personal-best or p-best, denoted as . The overall velocity at step can be expressed as:
where and are uniformly distributed random numbers between and . The constant weights and control the degree of convergence and exploration of the swarm. A rough set of values for the constants is (, , ) = (0.7, 0.35, 0.4), but fine tuning is possible depending upon the search. The positions of the particles are updated using the simple rule given below,
As an initial test of robustness of the PSO algorithm, we demonstrate the recovery of simulated BNS sources from a small population of about 500 injections in Gaussian noise. These are non-eccentric BNS injections with moderate aligned-spin components; the component masses are drawn from . To locate the optimal template parameters for the HL detector network, we used 5k particles and allowed up to 15 steps, aggregated to about 75k template evaluations and that many matched-filter computations per detector. This recovery corresponds to an aligned-spin search using PSO. A summary of simulation results are shown in the Fig. 2. The injected coincident SNR is the coincident SNR obtained from the template with the injected parameters. Here an average loss of about 0.4% is observed in the recovered coincident SNR from that of the injected values. We also observed that to tackle less dense regions of the search space, like that of the NSBH and BBHs, requires even lesser sampling of the parameter space in general. The final g-best location in the parameter space represents an optimal solution with the obtained template parameters . Here the point estimates of the parameters describing the simulated sources are reasonably unbiased, notably the chirp mass, and the effective spin parameter, as shown in Fig. 2. The modifications to this algorithm to adapt to the real instrumental noise are described in the next section.
II.3 Multiple detector optimization
When searching in the coincident mode, the processes described earlier in section II.2 can, in principle, proceed independently and in parallel for all the available detectors. However, in such a case, the templates are also computed independently for each detector and the point of convergence can differ due to several reasons. When projected from arbirary directions in the sky, the antenna pattern functions of the detetors modulate the incoming signal differently, thus effectively shifting the expected optimal point. In addition, the noise contributions from the individual detectors can slightly shift the convergence point. Otherwise, several local peaks form an elongated band of almost degenerate templates with similar SNRs as shown in the Fig. 3, which can confuse the convergence. But a true gravitational wave signal should have the times of arrival and the source parameters consistent across all the available detectors. To ensure the second criterion, the same set of template parameters should optimize the given fitness function. Besides, for a given template, a stretch of data can contain several noise triggers, however, such triggers are expected to be uncorrelated and appear at just one detector. Thus, to avoid spurious optimization of glitch triggers, a natural choice is to jointly optimize the reweighted SNRs across the network of detectors. The joint fitness function for a network of detectors is given by,
where is the reweighted SNR of the most likely trigger arising from an astrophysical signal at the th detector. Template points that trigger to a loud glitch in one of the detectors are downranked by the veto processes and the swarm further inhibits convergence to such template points over the detector network. However, at this stage, we do not conduct a time coincidence test of the triggers and is performed after the optimization. An additional advantage of the above procedure is that a single set of computation of the template waveforms serves the entire network of detectors. It is especially useful when generating the template waveforms is expensive as is generally the case in an eccentric search. A similar strategy is also employed to optimize the network SNR of the candidates found in the PyCBC Live search (Canton et al., 2021).
To evaluate the fitness function at any given iteration, the template waveform at is constructed and cross-correlated with the given stretch of data. Each of these results into an SNR timeseries whose ends can be corrupted due to filtering artifacts and are discarded. Triggers from the remaining portion of the SNR timeseries above a chosen threshold, typically , are examined for all the particles. The reweighted SNR values are computed at the trigger points using the power -test to reweigh the triggers. The reweighted SNR is designed to downrank the noise triggers while preserving the SNR values of signals modelled by templates. Additionally, we compute the ratio of the reweighted-SNR and the SNR for the triggers and use it to rank them for their astrophysical origin. Triggers with a ratio close to 1, typically above are retained and only the highest ranked trigger from each detector proceed for joint optimization over the network as given in (Eq. 4). Note that these primary triggers do not undergo any time coincidence test and are only used to arrive at the best guess of the g-best template over the detector network. Thus we proceed with the optimization to get the final g-best parameters and store the resulting triggers indefinitely. The coincidences are formed only using these secondary triggers from the final g-best template in a post-processing step. This is in contrast to the template-bank searches where each template in the bank can give rise to a foreground candidate and hence, also participate in the background estimation. However, it is still consistent that the triggers which are tested for foreground coincidences are used to estimate the background.
After updating the position-velocity, the algorithm also checks whether any particle has drifted beyond the desired search boundaries- a reflecting boundary puts back an outlier inside in a symmetric position while flipping its velocity. All particles undergo this condition after the newly proposed template positions using Eq. (3) is computed. While this is suitable in most of the cases, we find that an adsorbing, boundary is more effective for the eccentricity dimension towards the recovery of non-eccentric systems where instead of putting the outlier back to a symmetric location, we put it on the boundary. At this point, the current iteration is complete. The algorithm additionally checks for a terminal condition such as a convergence below a small mean radius of the swarm or a preset timeout. The boundary conditions can improve the termination and also ensures no hang-up occurs due to non-availability of template parameters outside the region allowed by a waveform model. Note that these steps, namely evolving the position-velocity, applying the boundary conditions and the termination checks are computationally trivial. A summary of the steps are shown in the flow chart in Fig. 4.
II.4 Significance of candidates
When the iterations finish, the algorithm returns the final g-best template point with the associated triggers from individual detectors. The g-best template is already consistent in source parameters over all the detectors. Now these triggers are tested for any zero-lag time coincidence to create foreground events. Several such foreground candidates, which pass both the parameter and the time coincidence tests, may still not be of astrophysical origin. To this end, one can assign a statistical significance to a foreground candidate and/or a probability of being of astrophysical origin, (Farr et al., 2015; Capano et al., 2017). Here we assign a false alarm rate, FAR to the foreground events using the widely used method of time-shifting of triggers obtained from processing the data immediately available before the candidate. If triggers obtained from a pair of detectors are artificially shifted with respect to each other by a time more than travel time of light between the two detectors, then the resulting coincidences cannot be of astrophysical origin. Such artificial background events are also ranked by their artificial network statistic value and a FAR estimate is assigned to a foreground candidate as
Here is the number of background events with a network statistic equal to or greater than the foreground event. is computed as the artificial time over which the background coincident events are manufactured. In the simple case of a two detector network, it can be approximated by , where is the coincident live time of the two detectors and is a non-astrophysical time-slide duration. To compute a multi-detector background, the triggers from the most sensitive detector may be slided with remaining detectors fixed (Davies et al., 2020). Even considering triggers from Virgo, a time-slide duration of 50 ms can be safely chosen and as many artificial events are created as possible. When additional detectors are ready with similar sensitivities, a general approach with each detector sliding may be appropriate. Thus, to attempt a significant confidence of FAR less than 1 per 100 years to a given foreground event, a buffer of a sufficient duration of recent triggers must be kept which turns out to be a minimum of 3.5 hrs in our case. Normally, a large window for buffering triggers is more exposed to less recent and abrupt changes in the detector noise. Online LVK searches, however, require a higher FAR threshold for considering candidates for follow-up processes. In this work, we have mostly used only 4096 seconds to reach FARs upto 1 per 10 years.
III Eccentric Search
In the previous section, we described a general framework to search for CBC signals using the PSO algorithm. Now, we demonstrate the algorithm specifically designed for spin-aligned eccentric sources. In the absence of the occasional and loud disturbances, or glitches, the real instrument data can be considered as nearly Gaussian. Thus in this section, we first demonstrate the recovery of simulated signals with our algorithm in coloured Gaussian noise. The signals are projected from random sky locations onto a network of simulated detectors and then injected into the noise segments. Here we at most use the 3-detector combination, from H1, L1, and V1, and use the publicly available O4 representative PSDs to colour the noise segments (https://dcc.ligo.org/LIGO T2000012/public, ). We then discuss the response of the search to the real data from the instruments when no likely signals are present. With each of these, we also include the corresponding observations for the aligned-spin searches. To reduce the cost of matched-filtering, data are downsampled to 1024 Hz while remaining sensitive to almost the full inspiral part of the signal as shown in the Fig. 5. Note that the TaylorF2Ecc is an inspiral-only model and does not anyway contain much power in the high frequencies. The loss in the SNR from such a high frequency cutoff would be just around 0.2 for a binary and even smaller for heavier binaries. We currently set the low frequency cutoff at 20 Hz. Since, the effects of eccentricity are more in the lower frequencies, it is reasonable to expect improvements with a lower low frequency cutoff.
III.1 Eccentric BNSs
In this section, we generate mock signals from a simulated population of about 500 eccentric BNS sources using the TEOBResumS-DALI model and inject them into the Gaussian noise simulated for the two LIGO detectors. Agnostic about the true astrophysical distribution, the merger component masses are uniformly drawn from , with moderately large values of the aligned-spin components in and small eccentricities from . We use TaylorF2Ecc to recover the signals using the PSO algorithm. The TaylorF2Ecc waveform model basically draws the parameter space coverage from the widely used TaylorF2 model to search for BNS systems. For long duration signals, we expect to see an insignificant SNR loss due to the missing merger-ringdown parts in the template waveforms as shown in (Fig. 6). We test the accuracy of the PSO algorithm to mainly recover the coincident SNR and the injected binary parameters. Note that our primary goal here is to demonstrate the effectiveness of the PSO algorithm, but it also tests the agreement of the waveform parameters between the two models. The choice of the injection distribution may not be astrophysically well informed, however, the luminosity distances to the simulated sources are chosen from Mpc distributed uniformly in volume, giving a realistic range of network SNRs for O4 and beyond. In Fig. 8, the accuracy in recovering the injected parameters is plotted for the component masses, the effective spin parameters and the eccentricity. We find that the PSO algorithm is readily sensitive to these small eccentrities with the total 75k matched-filtering operations per detector. Note that, even a small eccentricity can modulate the signal over a large number of cycles, attributing to the improvement in the recovered coincident SNR as shown in the Fig. 6. It also may be concluded that the waveform parameter values are largely precise for the two models from different formalisms. Note that whatever deviation observed here also includes the fact that the templates do not span the merger-ringdown phases of the signals.
III.2 Eccentric BBHs
Here we use the TEOBResumS-DALI waveform model for both the signal injection and the template generation. This waveform model contains the inspiral-merger-ringdown phases required for the SNR build up, especially for the high mass binaries. As before, Fig. 9 summarizes the recovered parameters for the H1 and L1 coincidences out of 1000 injections in the mass range , aligned-spins within a range of and eccentricities up to . The luminosity distances of the sources are chosen uniformly in volume between 500-3000 Mpc. The search was performed using about 30k matched-filtering operations per detector.
To further demonstrate capabilities of the eccentric search, we perform a simulation with a large population of about 10k BBH injections. We prepare two injection sets with identical source parameters except eccentricity as follows- (1) non-eccentric aligned-spin BBH injections using IMRPhenomPv2, and (2) eccentric aligned-spin BBHs using TEOBResumS-DALI. In this investigation, we use the HLV detector network with the projected O4 model sensitivities for the Advanced LIGO and the Advanced Virgo.
We simultaneously conduct the following two searches on both of the injection sets- both having identical configurations but (1) with eccentric spin-aligned templates using the TEOBResumS-DALI model, and (2) with non-eccentric spin-aligned templates using the IMRPhenomPv2 model. Since, the data segments are synthetically gaussian and hence do not contain noise glitches, we consider injections as found, once the triggers cross an SNR threshold of 5 in coincidence with at least two detectors. The search results are summarized in Fig. 10. We quickly note the following about the eccentric search.
The recovery rate is notably greater- an overall of more coincident events are recovered than that of the aligned-spin search.
The overall quality of recovery is better (e.g. the network SNRs) in comparison with the aligned-spin recovery for the commonly found injections (however, not explicit from the figure).
A fraction of extra events feature in the triple (HLV) coincidences with the eccentric search, indicating potential improvements their significance and the follow-ups.
Both the searches recover similar number and fractions of the total coincidences in the case of spin-aligned non-eccentric injections.
Note that these statistics can vary depending on the choice of the injection population, detector PSDs and the instrument noise with glitches playing a key role in the false alarm rate associated with an event. We discuss some of these challenges in the next section.
III.3 Background event rates
In the previous sections, we have shown that the eccentric searches can outperform the aligned-spin searches in the presence of Gaussian noise. However, the data from a real detector can contain nearby noise glitches which can overwhelm the searches. To investigate this, we start with an hypothesis that there is no strong astrophysical signal and data segments contain only detector noise and glitches. We perform the searches independently using same configuration as described in section III on several data chunks from the third observing run (O3a) of the Advanced LIGOs not containing any obvious CBC candidates (Abbott et al., 2023, 2021e). For simplicity we use data from only two detectors LHO and LLO. Note that the triggers crossing an SNR threshold undergo the same signal consitency tests and the network reweighted-SNR is used as the ranking statistic. For comparison, we also conduct the usual aligned-spin search on the same data segments with identical set-up. The results are summarized in the Fig. 11.
While searching for the BNS systems, we find that both the searches generate a similar rate of background events. The power -test is effective in this low mass region with long inspirals and vetoes out most of the glitches. On the other hand, the eccentric BBH search uses a time domain model for the templates, where the filter artefacts can lead to higher rate of background events if not conditioned properly. We constrain the waveform with the tapering scheme for time-domain templates illustrated in (McKechan et al., 2010). In addition, it is known that some of the glitches, such as the blips, can successfully dodge the power -test for the higher-mass templates. In the BBH searches, the component masses can go beyond and eccentricity can further shorten the template durations. Sine-gaussians can approximately model some of these glitches and separate them from GW signals by taking into account of the power at frequencies higher than those in the templates (Nitz, 2018). To benefit and extend the eccentric BBH search for high mass systems, this can be incorporated in the future when we expand the search space coverage.
III.4 Eccentric re-searching of GWTC events
Event |
Triggered |
GPS start |
FAR |
PSO |
PyCBC |
MBTA |
GstLAL |
GW150914 |
HL |
1126259461.8 |
46.1 |
26.3 |
0.002 |
0.073 |
0.38 |
23.7 |
23.7 |
23.6 |
24.4 |
GW151012 |
HL |
1128678898.5 |
26.6 |
17.6 |
0.099 |
0.283 |
0.38 |
9.7 |
9.8 |
9.5 |
10 |
GW170104 |
HL |
1167559935.7 |
40.4 |
20.7 |
-0.034 |
0.035 |
0.38 |
13.1 |
13.2 |
13 |
13 |
GW170729 |
HL |
1185389806.8 |
70.1 |
29.1 |
0.051 |
0.325 |
2.26 |
9.5 |
9.9 |
9.8 |
10.8 |
GW170809 |
HL |
1186302518.8 |
34.7 |
27.2 |
-0.186 |
0.277 |
0.38 |
10.2 |
10.8 |
12.2 |
12.4 |
GW170814 |
HL |
1186741860.5 |
37.5 |
26.9 |
0.179 |
0.2 |
0.38 |
12.7 |
15.0 |
16.3 |
15.9 |
GW170817 |
HL |
1187008882.4 |
1.66 |
1.15 |
-0.001 |
0.01 |
0.38 |
32.4 |
32.5 |
30.9 |
33.0 |
GW170818 |
HL |
1187058326.3 |
44.8 |
22.1 |
-0.185 |
0.218 |
0.38 |
9.9 |
10.0 |
11.3 |
GW170823 |
HL |
1187529256.1 |
51.2 |
43.8 |
0.264 |
0.056 |
0.38 |
11.3 |
11.6 |
11.1 |
11.5 |
GW190403_051519 |
GW190408_181802 |
HL |
1238782699.2 |
41.9 |
17.2 |
0.002 |
0.104 |
0.38 |
14.1 |
14.1 |
13.1 |
14.4 |
14.7 |
GW190413_052954 |
HL |
1239168612.1 |
54.9 |
43.6 |
0.126 |
0.071 |
0.38 |
8.7 |
8.7 |
8.5 |
GW190413_134308 |
HL |
1239198205.2 |
82.6 |
37.6 |
-0.297 |
0 |
0.38 |
9.2 |
9.5 |
9.3 |
10.3 |
10.1 |
GW190421_213856 |
HL |
1239917954.0 |
63.5 |
40.8 |
-0.148 |
0.108 |
0.38 |
10.2 |
10.2 |
10.1 |
9.7 |
10.5 |
GW190503_185404 |
HL |
1240944861.8 |
49.4 |
34.6 |
-0.097 |
0.204 |
0.38 |
12.2 |
12.2 |
12.2 |
12.8 |
12 |
GW190512_180714 |
HL |
1241719650.7 |
27.5 |
15.9 |
-0.068 |
0.12 |
0.38 |
11.4 |
11.6 |
12.4 |
11.7 |
12.2 |
GW190513_205428 |
11.6 |
13 |
12.3 |
GW190514_065416 |
8.4 |
GW190517_055101 |
HL |
1242107479.1 |
54.4 |
24.9 |
0.365 |
0.306 |
0.38 |
9.5 |
9.5 |
10.4 |
10.8 |
11.3 |
GW190519_153544 |
HL |
1242315362.3 |
92.6 |
85.3 |
-0.499 |
0.191 |
2.63 |
11.3 |
11.3 |
13.2 |
13.7 |
12.4 |
GW190521_030229 |
HL |
1242442967.1 |
143.5 |
133.8 |
0.223 |
0.246 |
0.009 |
13.5 |
13.7 |
13.7 |
13 |
13.3 |
GW190521_074359 |
HL |
1242459857.1 |
51.8 |
43.2 |
0.162 |
0.005 |
0.38 |
24.2 |
24.4 |
24 |
22.2 |
24.4 |
GW190527_092055 |
HL |
1242984073.1 |
60.7 |
25.1 |
0.088 |
0.322 |
0.38 |
9.0 |
9.0 |
8.7 |
GW190602_175927 |
HL |
1243533584.9 |
93.1 |
72.0 |
0.036 |
0.195 |
0.38 |
12.2 |
12.2 |
11.9 |
12.6 |
12.3 |
GW190620_030421 |
10.9 |
GW190630_185205 |
15.2 |
GW190701_203306 |
HLV |
1246048404.3 |
69.9 |
56.3 |
-0.235 |
0.436 |
2.63 |
12.3 |
12.3 |
11.9 |
11.3 |
11.7 |
GW190706_222641 |
HL |
1246487219.0 |
90.3 |
48.7 |
0.148 |
0.497 |
0.38 |
13.0 |
13.0 |
11.7 |
11.9 |
12.5 |
GW190708_232457 |
13.1 |
GW190719_215514 |
HL |
1247608532.4 |
53.9 |
30.6 |
0.184 |
0.164 |
0.38 |
8.5 |
8.5 |
GW190727_060333 |
HL |
1248242631.6 |
53.3 |
48.5 |
0.151 |
0.007 |
0.38 |
11.7 |
11.7 |
11.4 |
12 |
12.1 |
GW190731_140936 |
HL |
1248617394.1 |
66.1 |
16.7 |
-0.356 |
0.09 |
0.38 |
8.6 |
8.6 |
9.1 |
8.5 |
GW190803_022701 |
HL |
1248834439.5 |
58.6 |
36.7 |
-0.119 |
0.292 |
5.64 |
8.2 |
8.5 |
8.7 |
9 |
9.1 |
GW190805_211137 |
HL |
1249074715.0 |
88.5 |
35.2 |
0.326 |
0.335 |
0.38 |
7.9 |
7.9 |
GW190828_063405 |
HL |
1251009263.1 |
39.6 |
36.1 |
0.169 |
0.164 |
0.38 |
15.7 |
15.7 |
13.9 |
15.2 |
16.3 |
GW190828_065509 |
HL |
1251010526.0 |
29.7 |
13.9 |
0.075 |
0.158 |
0.75 |
11.0 |
11.0 |
10.5 |
11.1 |
10.8 |
GW190910_112807 |
13.4 |
GW190915_235702 |
HL |
1252627040.0 |
42.2 |
28.5 |
-0.116 |
0.175 |
0.38 |
12.8 |
12.8 |
13 |
12.7 |
13 |
GW190916_200658 |
HL |
1252699636.4 |
48.9 |
35.9 |
0.05 |
0.259 |
381.12 |
6.7 |
6.8 |
8.2 |
8.2 |
GW190925_232845 |
HV |
1253489341.3 |
28.2 |
14.8 |
0 |
0.127 |
0.38 |
9.9 |
9.9 |
9 |
9.4 |
GW190926_050336 |
HL |
1253509433.6 |
87.5 |
16.2 |
-0.208 |
0.338 |
1.13 |
8.1 |
8.1 |
9 |
GW190929_012149 |
HL |
1253755327.3 |
88.3 |
46.9 |
-0.062 |
0.14 |
0.75 |
9.6 |
9.6 |
9.4 |
10.3 |
10.1 |
GW191109_010717 |
HL |
1257296855.0 |
80.4 |
59.6 |
0.176 |
0.333 |
0.38 |
16.0 |
16.0 |
13.2 |
15.2 |
15.8 |
GW191127_050227 |
HL |
1258866165.0 |
46.7 |
26.5 |
-0.214 |
0 |
1.50 |
8.7 |
8.7 |
9.5 |
9.8 |
10.3 |
GW191204_110529 |
HL |
1259492746.7 |
35.5 |
27.6 |
0.131 |
0.023 |
0.38 |
9.6 |
9.6 |
8.9 |
8.1 |
9 |
GW191215_223052 |
HL |
1260484269.4 |
39.5 |
21.1 |
0.048 |
0.085 |
0.38 |
10.6 |
10.6 |
10.3 |
10.8 |
10.9 |
GW191222_033537 |
HL |
1261020954.9 |
68.5 |
51.6 |
0.018 |
0.092 |
0.38 |
11.4 |
11.8 |
11.5 |
10.8 |
12 |
GW191230_180458 |
HL |
1261764316.2 |
90.4 |
37.9 |
0.075 |
0.109 |
0.38 |
9.8 |
9.8 |
9.6 |
9.8 |
10.3 |
GW200112_155838 |
17.6 |
GW200128_022011 |
HL |
1264213229.6 |
70.6 |
52.2 |
0.25 |
0.078 |
0.38 |
10.1 |
10.1 |
9.8 |
9.4 |
10.1 |
GW200129_065458 |
HLV |
1264316115.7 |
40.4 |
31.4 |
0.156 |
0.201 |
0.38 |
20.8 |
25.5 |
16.3 |
26.5 |
GW200208_130117 |
HLV |
1265202095.5 |
51.7 |
33.5 |
-0.13 |
0.036 |
0.38 |
10.8 |
10.8 |
9.6 |
10.4 |
10.7 |
GW200208_222617 |
8.9 |
8.2 |
GW200209_085452 |
HL |
1265273709.9 |
67.8 |
37.1 |
-0.137 |
0.149 |
0.38 |
9.0 |
9.0 |
9.2 |
9.7 |
10 |
GW200216_220804 |
HL |
1265926102.6 |
79.5 |
32.7 |
0.107 |
0.12 |
0.38 |
8.8 |
8.8 |
9 |
8.8 |
9.4 |
GW200219_094415 |
HL |
1266140671.8 |
63.7 |
39.0 |
-0.042 |
0 |
1.13 |
10.0 |
10.0 |
9.9 |
10.6 |
10.7 |
GW200220_061928 |
GW200220_124850 |
8.2 |
8.2 |
GW200224_222234 |
HLV |
1266618171.9 |
49.6 |
36.8 |
0.105 |
0.169 |
0.38 |
14.8 |
18.1 |
19.2 |
19 |
18.9 |
GW200225_060421 |
HL |
1266645877.7 |
23.0 |
18.7 |
-0.037 |
0.088 |
0.38 |
11.1 |
12.3 |
12.3 |
12.5 |
12.9 |
GW200302_015811 |
10.6 |
GW200306_093714 |
7.8 |
8.5 |
GW200308_173609 |
7.9 |
8.3 |
8.1 |
GW200311_115853 |
HLV |
1267963150.8 |
40.7 |
32.2 |
-0.019 |
0.046 |
0.38 |
15.8 |
17.1 |
17 |
16.5 |
17.7 |
GW200322_091133 |
8 |
9 |
In this section, we re-analyze the GWTC BBH events with component masses lying between 10 and 100 available in the GWOSC (Abbott et al., 2023, 2021e). A general idea behind conducting such a re-analysis is to test whether the eccentric search can produce candidates in agreement with the earlier searches. This would enable us to test our capabilities for possibly contributing to conducting follow-up studies. We find that the eccentric search recovers a majority of the confident events with either three or a two detector coincidence without much fine-tuning. Also, note that however, the false alarm rate estimates here are limited by the duration of the analyzed data over which backgrounds are estimated. The backgrounds evaluated over longer durations of data would result into better FAR estimates in general.
To illustrate this, in the case of GW190521, we further extend the search space for the component masses to 150 and use past 3.5 hours of data to compute the false alarm rate as shown in the Fig. 12. To tackle a larger search parameter space coverage, the swarm size is increased to about 2400 particles while the number of steps are kept fixed. The resulting search recovers the event with a false alarm rate close to 1 per century.
Note that we have not considered any single detector events in our analyses, which we can address in the future. In table 1, the GPS start times correspond to the average times-of-arrival across the detectors. Unlike the merger time, like in the case of GW170817 where we have used a frequency domain template, which is fixed for an event, GPS start time provided here is the start time of the triggering template and can vary depending on the duration of the g-best template corresponding to the low frequency cutoff.
III.5 Runtime
Searches involving at least one neutron-star component are time-critical. To expedite the electromagnetic follow-up efforts, the LVK searches currently take a couple of seconds to few minutes to identify the event candidates. In the case of the template bank based searches, the process of matched-filtering can entirely proceed in parallel and the waveforms can be pre-computed and stored in the memory (Nitz et al., 2018). However, the speed of the PSO based searches is limited by the cost of computing templates on the fly and the serial iterations. Thus, we try to minimize the time spent at each iteration with task parallelism. We split the set of CPU intensive computations in a given iteration across the available CPU cores. Each share of the total computation is tackled by an independent process. The processes exchange the required information for computing the fitness and for communicating the results using Message Passing Interface (MPI). In other words, every core processes a fixed number of particles for all the interations. Each iteration is synchronized and a master process coordinates and drives the algorithm. The computations such as storing triggers to the disk for/or post-processing are also performed by the master process. Online analyses require that data be processed at least at the same rate as are recorded by the detectors. We have broadly timed the BNS search with 15 steps and about 5000 particles for the HL and the HLV networks as shown in Fig. 13. We observe that the runtime is almost inversely proportional to the number of CPU cores available.
IV Conclusions
We have explored a simple algorithm that allows us to search for compact binary coalescences without using a pre-computed template bank. Instead, the PSO algorithm chooses template points on the fly while exploring an arbitrary dimensional search space. We exploit this feature to implement a five dimensional search to include the orbital eccentricity, in addition to the standard search parameters, namely the component masses and the aligned spins. If astrophysical sources carry a non-negligible eccentricity, the eccentric search improves the recovered signal-to-noise ratios. In this work, we investigate a search for BNS systems using around 75k matched-filtering operations whereas a significant part of the BBH parameter space can be tackled with about 30k matched-filtering operations. The algorithm utilizes these many number of operations to iteratively arrive at an optimal detection statistic and also provides reasonably unbiased point estimates for the source parameters like the chirp mass, the effective spin parameter and the eccentricity. We have described our eccentric search pipeline for multiple detectors and demonstrated it on many GWTC events.
We have presented a parallelization scheme to lower the real-runtime of the search by utilizing multiple CPU cores. In the case of an eccentric search, the speed and the accuracy of the available waveform families, is of a general concern. Some are fast but do not support large portions of the plausible parameter space, for example in the large mass ratio limit for binaries like the NSBHs, while others become slow while covering all the regions. Searches including the effects due to spin precession coupled with eccentricity is an interesting avenue whenever accurate waveform models for the complex physics come about. The mean anomaly parameter can also be incorporated into the analyses when suitable waveforms are available. Fortunately, waveform physics is an active area of research and development. With efficient waveforms spanning larger parts of the parameter space and some fine-tuning of the proposed method, it is possible to deploy an eccentric search that can be used in low-latency for the current generation of ground based detectors.
V Acknowledgements
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
Most of our analyses use the existing tools in the PyCBC software package. We fruitfully benefited from the discussions in the LVK Collaboration, the LIGO India Scientific Collaboration (LISC) and the PyCBC group. We acknowledge the use of the Sarathi computing cluster at IUCAA for computational/numerical work. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The article reserves the LIGO Document number LIGO-P2300164. The research was carried out at the Center of Excellence in Space Sciences, India (CESSI). CESSI is a multi-institutional Center of Excellence hosted by the Indian Institute of Science Education and Research (IISER) Kolkata and has been established through funding from the Ministry of Education, Government of India. Support also comes from the Council for Scientific and Industrial Research (CSIR), India through File No:09/921(0272)/2019-EMR-I. We thank the anonymous referee for providing the valuable feedback on the manuscript.